# Decay-protected superconducting qubit with fast control enabled by integrated on-chip filters

### v1

@@ 0.06 $\mu$s
    Supplementary Fig. S7: Energy decay time $T_{1}$ of each qubit at different frequencies. a Measured $T_{1}$ of qubit Q1 (markers) and corresponding results of fitted models (dashed lines), Purcell decay (red) and dielectric-loss (yellow), as functions of qubit frequency. From the fit to the dielectric-loss model, we obtain a loss tangent of $3\times 10^{-6}$, manifesting as the dominant loss mechanism in our devices far from the resonance of the qubit and its readout resonator. b,c Measured $T_{1}$ of qubit (b) Q2 and (c) Q3 with an additional off-chip low-pass filter with a cut-off frequency at around 3 GHz installed in qubit control line.
    Furthermore, following a similar approach to the one used in Ref. [56], we show that the use of off-chip low-pass filter in the qubit control line, facilitating a passband at sub-harmonic frequency while introducing a stopband at qubit frequency, creates an unwanted resonance in the qubit drive line that modulates the measured $T_{1}$ in extreme ranges, as depicted in Supplementary Fig. S7b, c. By employing the integrated on-chip filters, we are not faced with such an issue and the resulting engineering and scaling challenge.

### v2

@@ Acknowledgments
Supplementary Fig. S7: Energy decay time $T_{1}$ of each qubit at different frequencies. a Measured $T_{1}$ of qubit Q1 (markers) and corresponding results of fitted models (dashed lines), Purcell decay (red) and dielectric-loss (yellow), as functions of qubit frequency. From the fit to the dielectric-loss model, we obtain a loss tangent of $3\times 10^{-6}$, manifesting as the dominant loss mechanism in our devices far from the resonance of the qubit and its readout resonator. b,c Measured $T_{1}$ of qubit (b) Q2 and (c) Q3 with an additional off-chip low-pass filter with a cut-off frequency at around 3 GHz installed in qubit control line. The error bars represent $1\sigma$ fitting uncertainty.
@@
Furthermore, following a similar approach to the one used in Ref. [60], we show that the use of an off-chip low-pass filter in the qubit control line, facilitating a passband at the subharmonic frequency while introducing a stopband at the qubit frequency, creates unwanted resonances in the qubit drive line that modulate the measured $T_{1}$ in extreme ranges, as depicted in Supplementary Fig. S7b, c. By employing the integrated on-chip filters, we are not faced with such an issue and the resulting engineering and scaling challenge.